OpenStudy (anonymous):

Write the equation of the line passing through points (3, -2) and (6, 4).

hartnn (hartnn):

first find the slope using, The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

OpenStudy (whpalmer4):

Once you have the slope, you can write an equation for the line of slope \(m\) passing through point \((x_0,y_0)\) with the point-slope formula \[y-y_0 = m(x-x_0)\]and then rearrange into a different form if desired.